# Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilites of Large Language Models via Game Play

### v2

@@
Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilites of Large Language Models via Game Play
@@ 5.3.1 Temporal Adaptation Across Games (Radar Overview)
Figure 1 summarises temporal reasoning patterns for llm_litellm_groq_llama3_70b_8192 across seven games. Each polygon corresponds to a game; axes are reasoning types and radii denote normalised frequencies over the course of play. We observe opponent-focused behaviour in tic_tac_toe and prisoners_dilemma, positional emphasis in connect_four, rule-based regularity in matching_pennies, and a larger uncategorised share in matrix_pd.
Figure 1: Temporal reasoning-type patterns across games for llm_litellm_groq_llama3_70b_8192.
@@ 5.3.2 Aggregate Per-Game Profiles (Stacked Bars and By-Game Views)
Figure 2: Stacked reasoning percentages by game: llm_litellm_groq_llama3_70b_8192.
Figure 3: Stacked reasoning percentages by game: llm_litellm_groq_llama3_8b_8192.
Figure 4: By-game reasoning percentages: Meta LLaMA 3.1 8B Instruct Turbo.
Figure 5: By-game reasoning percentages: llm_litellm_groq_llama3_70b_8192.
Figure 6: By-game reasoning percentages: llm_litellm_groq_llama3_8b_8192.
@@ 5.3.3 Within-Game Evolution (Same Game, Different Turns)
Figure 7: Within-game reasoning heatmap (tic_tac_toe): llm_litellm_groq_llama_3_1_8b_instant.
Figure 8: Turn-binned proportions (tic_tac_toe): llm_litellm_groq_llama_3_1_8b_instant.
Figure 9: Within-game reasoning heatmap (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
Figure 10: Turn-binned proportions (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
@@ 5.3.4 Per-Game Pie Charts (Detailed Breakdown Examples)
Figure 11: Pie breakdown (tic_tac_toe): llm_litellm_groq_llama_3.1_8b_instant.
Figure 12: Pie breakdown (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
Figure 13: Pie breakdown (kuhn_poker): llm_litellm_groq_llama3_70b_8192.
@@ 5.3.5 Entropy Across Agents Within a Game
Figure 14 reports entropy across agents for tic_tac_toe. The llm_litellm_groq_llama3_8b_8192 model shows a clear mid-game peak (greater strategic diversity), followed by a collapse to near-zero in late turns, consistent with convergence to Blocking. The other agents remain near zero throughout.

### v3

# Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilities of Large Language Models via Game Play

@@
Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilities of Large Language Models via Game Play
@@ 5.3.1 Temporal Adaptation Across Games (Radar Overview)
Figure 1 summarises temporal reasoning patterns for groq_llama3_70b_8192 across seven games. Each polygon corresponds to a game; axes are reasoning types and radii denote normalised frequencies over the course of play. We observe opponent-focused behaviour in tic_tac_toe and prisoners_dilemma, positional emphasis in connect_four, rule-based regularity in matching_pennies, and a larger uncategorised share in matrix_pd.
Figure 1: Temporal reasoning-type patterns across games for groq_llama3_70b_8192.
@@ 5.3.2 Aggregate Per-Game Profiles (Stacked Bars and By-Game Views)
Figure 2: Stacked reasoning percentages by game: groq_llama3_70b_8192.
Figure 3: Stacked reasoning percentages by game: groq_llama3_8b_8192.
Figures 4, 5, and 6 show that stronger models adapt their reasoning to the structural demands of each game. For example, the Kimi-K2-Instruct model relies almost exclusively on opponent modeling in Kuhn Poker and Prisoner’s Dilemma, while exhibiting more diverse strategies in Connect Four and Tic-Tac-Toe. GPT-4, in turn, displays greater adaptability by deploying blocking strategies in Connect Four while relying on opponent modeling in sequential games. GPT-3.5-Turbo, however, exhibits weaker specialization, with many outputs uncategorized and lacking clear strategic grounding. These findings suggest that stronger models not only adapt within games but also calibrate their strategies across games.
Figure 4: Reasoning distribution across games for kimi_k2_instruct. Opponent modeling dominates in simpler games, while more complex board games elicit a broader range of reasoning types.
Figure 5: Reasoning distribution for gpt_4. The model adapts its reasoning to game structure, mixing opponent modeling, blocking, and uncategorized strategies.
Figure 6: Reasoning distribution for gpt_3.5_turbo. Compared to GPT-4, the model shows weaker adaptation and a higher proportion of uncategorized reasoning.
@@ 5.3.3 Within-Game Evolution (Same Game, Different Turns)
Figure 7: Within-game reasoning heatmap (tic_tac_toe): groq_llama_3_1_8b_instant.
Figure 8: Turn-binned proportions (tic_tac_toe): groq_llama_3_1_8b_instant.
Figure 9: Within-game reasoning heatmap (tic_tac_toe): groq_llama3_8b_8192.
Figure 10: Turn-binned proportions (tic_tac_toe): groq_llama3_8b_8192.
@@ 5.3.4 Per-Game Pie Charts (Detailed Breakdown Examples)
Figure 11: Pie breakdown (tic_tac_toe): groq_llama_3.1_8b_instant.
Figure 12: Pie breakdown (tic_tac_toe): groq_llama3_8b_8192.
Figure 13: Pie breakdown (kuhn_poker): groq_llama3_70b_8192.
@@ 5.3.5 Entropy Across Agents Within a Game
Figure 14 reports entropy across agents for tic_tac_toe. The groq_llama3_8b_8192 model shows a clear mid-game peak (greater strategic diversity), followed by a collapse to near-zero in late turns, consistent with convergence to Blocking. The other agents remain near zero throughout.